The cafe where you work just ran out of coffee, and you are at the store to buy 11

2


112 pounds of coffee. You have put a can with 3

4


34 pound of coffee into your shopping cart. How many more pounds of coffee do you need?



1
Expert's answer
2021-06-18T12:41:52-0400
"1\dfrac{1}{2}-\dfrac{3}{4}=\dfrac{3}{2}-\dfrac{3}{4}=\dfrac{6-3}{4}=\dfrac{3}{4}"

I need 3/4  pounds of coffee .